OpenStudy (anonymous):

182.31 = v ^ 0.16 how do i find v?

OpenStudy (anonymous):

\[182.31 = v^0.16\] \[182.31^{1 \over 0.16} = v^{0.16 \over 0.16} = v^1 = v \approx 1.349*10^{14} \] It's the same thing as with quadratic other exponents . If the exponent is 2 you take the square-root... in this case, you take the 0.16th-root.. \[182.31 ^ {1 \over 0.16}= \sqrt[0.16]{182.31}\]